OpenStudy (anonymous):

What is the equation of the circle passing through (-4, -4) and tangent to the line 2x – 3y + 9 = 0 at (-3, 1)?

OpenStudy (freckles):

Calculus? Algebra?

OpenStudy (anonymous):

Calculus

OpenStudy (freckles):

So we know we have a circle... \[(x-h)^2+(y-k)^2=r^2 \]

OpenStudy (freckles):

we need to differentiate this

OpenStudy (freckles):

we know to do this because we are given something about a tangent line to the circle so we are hoping to use that

OpenStudy (freckles):

2x-3y+9=0 -3y=-2x+9 y=2 x/ 3 - 3 slope is 2/3 you check me on that if you want... But can you differentiate the equation I gave you w.r.t to x h,k, and r are constants

OpenStudy (anonymous):

so what will i do if i found the slope of the tangent line? I don't get it.

OpenStudy (freckles):

y'=2/3 but you need to find y' first

OpenStudy (anonymous):

why y'? m=slope=2/3

OpenStudy (freckles):

the slope of tangent line at x=a is the derivative of the function/relation evaluated at x=a

OpenStudy (freckles):

by the way let me fix something 2x-3y+9=0 -3y=-2x-9 y=2x/3+3 slope is still 2/3 for this line we want to find y' and evaluate y' at the point (-3,1)

OpenStudy (freckles):

can you do that?

OpenStudy (anonymous):

i'll try to analyze

OpenStudy (freckles):

\[\frac{d}{dx}(x-h)^2=? \\ \frac{d}{dx}(y-k)^2=? \\ \frac{d}{dx}(r^2)=?\]

OpenStudy (freckles):

Can you compute these derivatives?

OpenStudy (freckles):

I will do the second one for you \[\frac{d}{dx}(y-k)^2=2(y-k)y'\]

OpenStudy (freckles):

the first one is very similar and you should know the last one is pretty easy because derivative of a constant is?

OpenStudy (anonymous):

ohhh our professor haven't teach us that yet

OpenStudy (anonymous):

in specific the lessons is analytic geometry

OpenStudy (anonymous):

but our class is calculus 1

OpenStudy (freckles):

ok but we aren't suppose to use calculus right? just algebra and geometry ? --- \[y=\frac{2}{3}x+3 \text{ is the tangent line of the circle at } (-3,1) \\ \text{ I want to determine the radius } \\ \text{ pretending the center is } (h,k) \\ \text{ We can draw a line from } (h,k) \text{ \to } (-3,1) \\ \text{ We know this line is perpendicular to the tangent line } \\ y=\frac{-3}{2}x+B \\ \text{ we can determine B by using our point } (-3,1) \\ 1=\frac{-3}{2}(-3)+B \\ 1=\frac{9}{2}+B \\ \text{ solve for B to find the y-intercept of this tangent line }\]